OpenStudy (anonymous):

Find the perimeter of a triangle with vertices (2,4) (-1,3) and (2, -3). (show your work on graph paper)

OpenStudy (amistre64):

we need 3 distance how do we determine distance between 2 points?

OpenStudy (amistre64):

A (2, 4) B(-1,3) C(2, -3) -2-4 -2-4 -2 -4 -------------------- 0,0 -3,-1 0,-7 AC = 7 by default AB = the square sum and sqrt of its parts = sqrt(9+1) = sqrt(10) now we just need the measure from B to C to add in

OpenStudy (amistre64):

B(-1,3) C(2,-3) 1 -3 1 -3 --------------- 0,0 3,0 by default, id say BC = 3

OpenStudy (amistre64):

our perimeter then should be: 7+3+sqrt(10)

OpenStudy (anonymous):

so the answer is 10+sqrt(10)?

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

can u show how to do it using graph paper?

OpenStudy (amistre64):

no, i cant lol

OpenStudy (amistre64):

you plot your points and determine the distance from one point to another using the same formulas

OpenStudy (amistre64):

or a ruler i spose

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you check on my other question...i need help lol